OpenStudy (anonymous):

Solve the equations and check your solutions. If there is no solution, say so 1/x + 2/3x= 1/3 need the step please. Thanks in advance. 5/(4x )+2=1/x 10/(3x-8)=2/x 4- 8x/(x+1)=8/(x+1) 3x/(x-2)=(3x+5)/(x-1) I have a hard time to solve this equations.

OpenStudy (amistre64):

multiply at all by 3x to clear the fractions

OpenStudy (amistre64):

i missed an x lol

OpenStudy (amistre64):

(3x)1/x + (3x)2x/3= (3x)1/3 (3)1/1 + (x)2x/1= (x)1/1 3 + 2x^2 = x